OpenStudy (rstew):

Given two functions, f1(x) = 2x and f2(x) = x2, explain what ∫20[f1(x)−f2(x)]dx represents and calculate this integral in the interval [0, 2]. Show all work

OpenStudy (sshayer):

\[\int\limits_{0}^{2}20\left[ f1(x)-f2(x) \right]dx=\int\limits_{0}^{2}20\left[ 2x-x^2 \right]dx=?\]

OpenStudy (dumbcow):

The integral represents the area between the graphs of 2x and x^2